Discuss possible chemical reactions from the mixing of the chemicals or their breakdown of Gasoline, Phosgene, and Oxygen Gas
1
Expert's answer
2013-01-24T08:32:49-0500
Phosgene reactswith oxygen from the equation:
2COCl2 + O2 = 2CO2 + 2Cl2
Phosgene mixed with gasoline in any proportions.
Gasoline consisting mainly of heptaneand octane reacts with oxygen as follows:
С7H16 + 11O2 = 7CO2 + 8H2O
2C8H18 + 25O2 = 16CO2 + 18H2O